# Measuring football fever through wearable technology: A case study on the German cup final

### v1

@@ Introduction
For fans, football is an emotional ride — full of joy, frustration, hope, and pride — whether they are cheering in the stands, watching from home, or keeping up with updates on their phones. This collective enthusiasm fuels a thriving football economy, from ticket sales and merchandise to media rights and sponsorship deals (Sauer et al.,, 2024). Unsurprisingly, the cultural and commercial significance has inspired a growing body of research that explores football as both an emotional experience and a social phenomenon (Ötting et al.,, 2023; Michels et al.,, 2023). This paper seeks to quantify physiological and emotional responses through a unique case study in German professional football, examining their intensity, variability, and underlying drivers. It draws on data from fan engagement during the 2025 cup final in Germany, which was arguably the most important match in the history of Arminia Bielefeld, one of the two participating clubs.
@@
The literature documents that fans often develop strong psychological bonds with their team, with concepts such as team identification and identity fusion describing how fans feel “at one” with their club (Newson et al.,, 2020). This deep attachment amplifies emotional reactions to match events, triggering substantial physiological responses. Research has recorded increases in spectators’ heart rate, blood pressure, and stress hormones during high-stakes matches (van der Meij et al.,, 2012). Fans who are firmly fused with their team experience the highest release of stress hormones when under game stress (Newson et al.,, 2020). In extreme cases, the cardiovascular strain of spectating can be severe, leading to an increased incidence of cardiac emergencies (Maturana et al.,, 2021; Mahečić et al.,, 2022). Notably, the communal rituals surrounding matches can evoke powerful shared emotions: For example, the excitement of Brazilian football fans (measured via heart rate arousal) has been shown to not peak during the match itself but during pre-match fan rituals, with only goal celebrations matching that pre-match high (Xygalatas et al.,, 2025). Summarising, this literature underscores how profoundly sports fandom engages both mind and body, providing a foundation for quantifying fans’ physical and emotional responses in our case study.
@@
The smartwatch records various sports-related metrics, including active seconds, motion intensity, and hourly steps. For this study, we primarily focus on the heart rate and stress level as indicators of fans’ emotional attachment to the match. This choice is consistent with previous research demonstrating the suitability of these variables for capturing emotional responses (Appelhans and Luecken,, 2006). The heart rate is reported as beats per minute (bpm) and recorded in 15-second intervals, whereas the stress level is recorded in three-minute intervals. The stress level tracked by the smartwatch is estimated from heart rate variability and expressed on a scale from 0 to 100, with 0 representing no stress and 100 indicating maximum stress. Notably, the stress level is not recorded when the person wearing the smartwatch is highly active (e.g. during exercise), as physical exertion itself induces substantial heart rate variability. Descriptive statistics on both stress levels and heart rates are provided in Table 1 and Figure 2 below.
@@
Given the established link between age and heart rate (Tanaka et al.,, 2001; Reardon and Malik,, 1996), we incorporated personal characteristics within category (1) of the survey. Beyond age and gender, we expect the degree of engagement with Arminia Bielefeld to influence fans’ physiological responses during the match. Research in basketball, for instance, has shown that in-person spectatorship produces greater group physiological synchrony and more intense “transformative” experiences than viewing in a small remote group (Baranowski-Pinto et al.,, 2022). In our case, a dedicated fan who regularly attends matches of Arminia Bielefeld in the stadium might exhibit a different physiological reaction to match dynamics compared with someone who rarely attends in person. We therefore asked for the attachment to Arminia Bielefeld, operationalised through club membership status and the number of matches attended in person during the 2024/25 season.
Under category (2) of the survey, we asked participants where they had followed the cup final. Those who reported watching the match (either on TV, at a public gathering, or in the stadium) were also asked whether they had consumed alcohol (Strüven et al.,, 2025). Category (3) applied exclusively to respondents who attended the final in the stadium. These participants were asked about their arrival time in the host city to capture potential travel-related impact on the vital parameters. Media reports indicated that overcrowding at stadium entrances caused delays, potentially evoking feelings of danger, anxiety, or frustration, which could in turn influence physiological parameters. Accordingly, we asked participants when they joined the entrance queue and how long the entry process lasted. In addition, participants were asked about betting behaviour related to the cup final; however, no one reported having placed a bet.
@@
Variable Time window Minimum 1. Quartile Median Average 3. Quartile Maximum
Stress level Cup final 21.0 28.3 48.1 45.3 58.2 70.5
Regular day 13.5 23.3 34.6 31.9 39.3 47.5
Overall 13.5 24.4 35.1 33.6 40.0 70.5
@@
Table 1: Summary statistics of stress levels and heart rates for individual participants, measured at single time points, on the day of the cup final (May 24, 2025) and on regular days within the observation period (May 19 to July 31, 2025).
@@
According to pre-match betting odds, provided by a major European bookmaker, VfB Stuttgart, the runner-up of the previous national league season and opponent of Arminia Bielefeld, was deemed the clear favourite. Specifically, translating these betting odds into odds-implied winning probabilities as a predictor for the match outcome (Winkelmann et al.,, 2024), the estimated probability of a VfB Stuttgart victory was 70%, a draw after 90 minutes was about 20%, and an Arminia Bielefeld win was approximately 10%. Although the initial phase of the match was quite balanced, with even promising scoring opportunities for Arminia Bielefeld, VfB Stuttgart scored in the 14th, 22nd, and 27th minutes. This 3-0 lead increased VfB Stuttgart’s odds-implied winning probability to over 95%, rising to more than 97% by halftime. In the second half, VfB Stuttgart scored again in the 67th minute, before Arminia Bielefeld was able to reduce the deficit with goals in the 84th and 87th minutes, eventually resulting in a 4–2 victory for VfB Stuttgart. Despite these being the first goals ever scored by a third-division club in the cup final, they had minimal impact on the odds-implied probabilities: Prior to Arminia Bielefeld’s first goal, the probabilities were 99.5% for a VfB Stuttgart win and 0.25% each for a draw and an Arminia Bielefeld victory. Following the 4–2 scoreline, VfB Stuttgart’s winning probability only slightly decreased to 97.7%, with a 2.0% probability of a draw and 0.3% for an Arminia Bielefeld win.
@@
To further contextualise the fan experience, Figure 5 shows the average heart rate over the course of the match, while distinguishing between participants who watched the match in the stadium, at a public gathering, or on TV. These values are based on different numbers of participants, determined by the location of the match attendance, and include only those participants who filled out the additional survey. The average heart rate is highest for participants in the stadium (average: 94.2 bpm), followed by those watching TV (79.4 bpm) and at a public gathering (73.8 bpm). On average, heart rates of stadium attendees are 23.1% higher than those of participants watching elsewhere. After the first goal of Arminia Bielefeld, this difference increases to a maximum of 35.8%, with a maximum average heart rate of 108.0 bpm in the stadium. These findings suggest that the emotional and physiological arousal associated with watching the match is particularly pronounced in the stadium environment compared with other viewing contexts. Beyond the location where the match was watched, alcohol consumption is also associated with elevated heart rates: The average heart rate of participants who reported alcohol intake is on average 5.3% higher throughout the match, 7.4% higher during the second half, and even 11.7% higher following Arminia Bielefeld’s first goal. Interpretation of the heart rate should therefore take into account that alcohol intake can contribute to an acute increase in cardiovascular strain, particularly in emotionally arousing contexts: Elevated heart rates in combination with alcohol are known to increase the risk of arrhythmias and other adverse cardiac events (Brunner et al.,, 2024), which may be of clinical relevance during high-stress situations such as sports events.
@@ Discussion
Still, limitations remain: reliance on indirect stress measures provided by the smartwatches, a small subsample of the post-study survey, and the focus on a single exceptional event. While our study relies on smartwatch heart rate and heart rate variability-based stress measures, future research could profit from integrating broader physiological variables. Recent advances in wearable technology illustrate the potential of multimodal monitoring. For instance, research demonstrated that multispectral sensor fusion in smartwatches allows continuous assessment of hydration and sweat loss, offering new perspectives on physiological reactions during emotionally intense events(Volkova et al.,, 2023). Similarly, hybrid wearable patches that combine electromyography with sweat cortisol sensing can capture both muscular and biochemical stress indicators(Hossain et al.,, 2024). Incorporating such multimodal data would enable a richer and more precise characterisation of football fever, contributing to a comprehensive understanding of fan arousal. Beyond the scope of the present work, our study opens avenues for examining vital parameters in daily life and within general natural high‑pressure contexts.

### v2

@@ Introduction
For fans, football is an emotional ride — full of joy, frustration, hope, and pride — whether they are cheering in the stands, watching from home, or keeping up with updates on their phones. This collective enthusiasm fuels a thriving football economy, from ticket sales and merchandise to media rights and sponsorship deals (Sauer et al., 2024). Unsurprisingly, the cultural and commercial significance has inspired a growing body of research that explores football as both an emotional experience and a social phenomenon (Ötting et al., 2023; Michels et al., 2023). This paper seeks to quantify physiological and emotional responses through a unique case study in German professional football, examining their intensity, variability, and underlying drivers. It draws on data from fan engagement during the 2025 cup final in Germany, which was arguably the most important match in the history of Arminia Bielefeld, one of the two participating clubs.
@@
The literature documents that fans often develop strong psychological bonds with their team, with concepts such as team identification and identity fusion describing how fans feel “at one” with their club (Newson et al., 2020). This deep attachment amplifies emotional reactions to match events, triggering substantial physiological responses. Research has recorded increases in spectators’ heart rate, blood pressure, and stress hormones during high-stakes matches (van der Meij et al., 2012). Fans who are firmly fused with their team experience the highest release of stress hormones when under game stress (Newson et al., 2020). In extreme cases, the cardiovascular strain of spectating can be severe, leading to an increased incidence of cardiac emergencies (Maturana et al., 2021; Mahečić et al., 2022). Notably, the communal rituals surrounding matches can evoke powerful shared emotions: For example, the excitement of Brazilian football fans (measured via heart rate arousal) has been shown to not peak during the match itself but during pre-match fan rituals, with only goal celebrations matching that pre-match high (Xygalatas et al., 2025). Summarising, this literature underscores how profoundly sports fandom engages both mind and body, providing a foundation for quantifying fans’ physical and emotional responses in our case study.
@@
The smartwatch records various sports-related metrics, including active seconds, motion intensity, and hourly steps. For this study, we primarily focus on the heart rate and stress level as indicators of fans’ emotional attachment to the match. This choice is consistent with previous research demonstrating the suitability of these variables for capturing emotional responses (Appelhans and Luecken, 2006). The heart rate is reported as beats per minute (bpm) and recorded in 15-second intervals, whereas the stress level is recorded in three-minute intervals. The stress level tracked by the smartwatch is estimated from heart rate variability and expressed on a scale from 0 to 100, with 0 representing no stress and 100 indicating maximum stress. Notably, the stress level is not recorded when the person wearing the smartwatch is highly active (e.g. during exercise), as physical exertion itself induces substantial heart rate variability. Descriptive statistics on both stress levels and heart rates are provided in Table 1 and Figure 2 below.
@@
Given the established link between age and heart rate (Tanaka et al., 2001; Reardon and Malik, 1996), we incorporated personal characteristics within category (1) of the survey. Beyond age and gender, we expect the degree of engagement with Arminia Bielefeld to influence fans’ physiological responses during the match. Research in basketball, for instance, has shown that in-person spectatorship produces greater group physiological synchrony and more intense “transformative” experiences than viewing in a small remote group (Baranowski-Pinto et al., 2022). In our case, a dedicated fan who regularly attends matches of Arminia Bielefeld in the stadium might exhibit a different physiological reaction to match dynamics compared with someone who rarely attends in person. We therefore asked for the attachment to Arminia Bielefeld, operationalised through club membership status and the number of matches attended in person during the 2024/25 season.
Under category (2) of the survey, we asked participants where they had followed the cup final. Those who reported watching the match (either on TV, at a public gathering, or in the stadium) were also asked whether they had consumed alcohol (Strüven et al., 2025). Category (3) applied exclusively to respondents who attended the final in the stadium. These participants were asked about their arrival time in the host city to capture potential travel-related impact on the vital parameters. Media reports indicated that overcrowding at stadium entrances caused delays, potentially evoking feelings of danger, anxiety, or frustration, which could in turn influence physiological parameters. Accordingly, we asked participants when they joined the entrance queue and how long the entry process lasted. In addition, participants were asked about betting behaviour related to the cup final; however, no one reported having placed a bet.
@@
Variable Time window Min. 1. Quartile Median Average 3. Quartile Max.
Stress level Cup final 0.0 21.0 40.0 44.2 67.0 100.0
Regular day 0.0 16.0 24.0 31.1 43.0 100.0
Overall 0.0 16.0 25.0 31.3 43.0 100.0
@@
Table 1: Summary statistics of stress levels (scale from 0 to 100) and heart rates (bpm) for individual participants, measured at single time points, on the day of the cup final (May 24, 2025) and on regular days within the observation period (May 14 to July 31, 2025).
@@
According to pre-match betting odds, provided by a major European bookmaker, VfB Stuttgart, the runner-up of the previous national league season and opponent of Arminia Bielefeld, was deemed the clear favourite. Specifically, translating these betting odds into odds-implied winning probabilities as a predictor for the match outcome (Winkelmann et al., 2024), the estimated probability of a VfB Stuttgart victory was 70%, a draw after 90 minutes was about 20%, and an Arminia Bielefeld win was approximately 10%. Although the initial phase of the match was quite balanced, with even promising scoring opportunities for Arminia Bielefeld, VfB Stuttgart scored in the 14th, 22nd, and 27th minutes. This 3-0 lead increased VfB Stuttgart’s odds-implied winning probability to over 95%, rising to more than 97% by halftime. In the second half, VfB Stuttgart scored again in the 67th minute, before Arminia Bielefeld was able to reduce the deficit with goals in the 84th and 87th minutes, eventually resulting in a 4–2 victory for VfB Stuttgart. Despite these being the first goals ever scored by a third-division club in the cup final, they had minimal impact on the odds-implied probabilities: Prior to Arminia Bielefeld’s first goal, the probabilities were 99.5% for a VfB Stuttgart win and 0.25% each for a draw and an Arminia Bielefeld victory. Following the 4–2 scoreline, VfB Stuttgart’s winning probability only slightly decreased to 97.7%, with a 2.0% probability of a draw and 0.3% for an Arminia Bielefeld win.
@@
To further contextualise the fan experience, Figure 5 shows the average heart rate over the course of the match, while distinguishing between participants who watched the match in the stadium, at a public gathering, or on TV. These values are based on different numbers of participants, determined by the location of the match attendance, and include only those participants who filled out the additional survey. The average heart rate is highest for participants in the stadium (average: 94.2 bpm), followed by those watching TV (79.4 bpm) and at a public gathering (73.8 bpm). On average, heart rates of stadium attendees are 23.1% higher than those of participants watching elsewhere. After the first goal of Arminia Bielefeld, this difference increases to a maximum of 35.8%, with a maximum average heart rate of 108.0 bpm in the stadium. These findings suggest that the emotional and physiological arousal associated with watching the match is particularly pronounced in the stadium environment compared with other viewing contexts. Beyond the location where the match was watched, alcohol consumption is also associated with elevated heart rates: The average heart rate of participants who reported alcohol intake is on average 5.3% higher throughout the match, 7.4% higher during the second half, and even 11.7% higher following Arminia Bielefeld’s first goal. Interpretation of the heart rate should therefore take into account that alcohol intake can contribute to an acute increase in cardiovascular strain, particularly in emotionally arousing contexts: Elevated heart rates in combination with alcohol are known to increase the risk of arrhythmias and other adverse cardiac events (Brunner et al., 2024), which may be of clinical relevance during high-stress situations such as sports events.
@@ Discussion
Still, limitations remain: reliance on indirect stress measures provided by the smartwatches, a small subsample of the post-study survey, and the focus on a single exceptional event. While our study relies on smartwatch heart rate and heart rate variability-based stress measures, future research could profit from integrating broader physiological variables. Recent advances in wearable technology illustrate the potential of multimodal monitoring. For instance, research demonstrated that multispectral sensor fusion in smartwatches allows continuous assessment of hydration and sweat loss, offering new perspectives on physiological reactions during emotionally intense events(Volkova et al., 2023). Similarly, hybrid wearable patches that combine electromyography with sweat cortisol sensing can capture both muscular and biochemical stress indicators(Hossain et al., 2024). Incorporating such multimodal data would enable a richer and more precise characterisation of football fever, contributing to a comprehensive understanding of fan arousal. Beyond the scope of the present work, our study opens avenues for examining vital parameters in daily life and within general natural high‑pressure contexts.